# Computational bottlenecks for denoising diffusions

### v1

@@ 3 Main results
For technical reasons, we will state two separate results for the probability distribution $\mu=\mu_{n,k}$ described in the introduction, depending on the scaling of $k$ with $n$: in Section 3.1 we assume $\sqrt{n}\ll k\ll n$; while in Section 3.3 we assume $k\ll\sqrt{n}$. The reason for this distinction is that the nature of the estimator $\hat{\boldsymbol{m}}$ changes at the threshold $k\asymp\sqrt{n}$.
@@
As discussed below, see Conjecture 3.2, it is expected that no polynomial-time algorithm can estimate $\boldsymbol{x}$ significantly better than the null estimator $\hat{\boldsymbol{m}}_{\mbox{\tiny\rm null}}=\boldsymbol{0}$ for ${\underline{k}}_{n}\leq k\ll n$, with ${\underline{k}}_{n}\ll n$ a suitable sequence.
We also note that since $\|\boldsymbol{x}\|_{F}=1$ almost surely under the target distribution $\boldsymbol{x}\sim\mu$, the Bayes optimal denoiser $\boldsymbol{m}(\,\cdot\,)$ does not depend explicitly on $t$ and we will therefore use the notation $\boldsymbol{m}(\boldsymbol{y}_{t})$.
@@ 3.1 Moderately sparse regime: $\sqrt{n}\ll k\ll n$
We next state our main result in the moderately sparse regime $\sqrt{n}\ll k\ll n$. Given an arbitrary polynomial time computable drift $\hat{\boldsymbol{m}}_{0}$, we construct a different poly-time drift $\hat{\boldsymbol{m}}$, with nearly equal score matching objective and yet incorrect sampling. In the next subsection we state a related result that establishes incorrect sampling from an explicit denoiser with near-optimal denoising properties.
@@
In order to connect this theorem with the general discussion in the introduction, we recall two facts from the literature on submatrix estimation: $(i)$ The ideal posterior mean estimator $\boldsymbol{m}(\boldsymbol{y}_{t})$ achieves small mean square error in a large interval below $t_{\mbox{\tiny\rm alg}}$ (Proposition 3.1); $(ii)$ No polynomial-time estimator is expected to perform substantially better than the null estimator below $t_{\mbox{\tiny\rm alg}}$ (Conjecture 3.2).
To provide more detail on point $(i)$, we state below a characterization of the Bayes optimal error. The proof is essentially the same (indeed simpler) as the main result in [BIS15], which considers the case of asymmetric matrices. An alternative proof (for $k\leq n^{a}$, $a<5/6$) also follows from [BMR20].
Proposition 3.1 (Modification of [BIS15]).
Let $\boldsymbol{m}(\,\cdot\,)$ be the posterior mean estimator in the Gaussian submatrix problem (8). Assume $1\ll k\ll n$, and define $t_{\mbox{\tiny\rm Bayes}}(n,k):=2k\log(n/k)$. Then, for any $\delta>0$, we have
@@
Several groups have proved hardness results for estimating $\boldsymbol{u}$ in the model considered here, for $t$ sufficiently smaller than the spectral threshold $n/2$. In particular [BBH18] establishes hardness conditional on hardness on the planted clique problem for $t\leq n^{1-\varepsilon}$, $\varepsilon>0$, [HKP+17] proved failure of sum-of-squares relaxations also for $t\leq n^{1-\varepsilon}$, [SW22] proved hardness for low-degree polynomials for the related hidden submatrix problem for $t\leq cn$, $c>0$ a sufficiently small constant. Finally, [KWB19] prove hardness for low-degree polynomials algorithms when $t\leq(1-\varepsilon)n/2$, albeit their results require $k$ to be of order $n$.
@@
Remark 3.1.
It makes sense to assume that $\|\hat{\boldsymbol{m}}_{0}(\cdot,\cdot)\|_{F}\leq 1$. Since $\boldsymbol{x}\sim\mu_{n,k}$ and $\text{supp}(\mu_{n,k})\subset\{\boldsymbol{a}:\|\boldsymbol{a}\|_{F}\leq 1\}$ and the latter is a convex set, projecting any estimator $\hat{\boldsymbol{m}}_{0}$ onto this set yields a smaller meansquare error.
@@ 3.2 Moderately sparse regime: A concrete example
7:  Compute $\hat{S}$ by
$\hat{S}:=\Big{\{}i\in[n]:\;|v_{t,i}|\geq\frac{\varepsilon}{\sqrt{k}}\Big{\}}\,;$ (13)
@@
    $\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C\,e^{-cn/k}\,.$
@@ 3.3 Very sparse regime: $k\ll\sqrt{n}$
3:  Let $\boldsymbol{g}_{t}\sim{\mathsf{N}}(0,t{\boldsymbol{I}}_{n\times n})$ and compute
$\boldsymbol{y}_{t,+}:=\boldsymbol{y}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t},,\;\;\;\;\boldsymbol{y}_{t,-}:=\boldsymbol{y}_{t}-\sqrt{\frac{1}{\varepsilon}}\boldsymbol{g}_{t}\,.$ (14)
@@
    $\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C(\delta,D)n^{-D}\,.$
@@
The pseudocode for the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$ that is constructed in the above is given as Algorithm 2. This is based on a standard approach in the literature [DM16, CLR17], with some modifications to allow for its analysis in the diffusion setting. The main steps are as follows:
  • Perform Gaussian data splitting of $\boldsymbol{y}_{t}$ into $\boldsymbol{y}_{t,+}$, $\boldsymbol{y}_{t,-}$, see Eq. (14), with most of the information preserved in $\boldsymbol{y}_{t,+}$.
  • Use entrywise soft thresholding $\eta_{s}(x)=(|x|-s)_{+}\operatorname{sign}(x)$ to reduce the noise in the symmetrized version of $\boldsymbol{y}_{t,+}$.
  • Compute a first estimate of the latent vector $\boldsymbol{1}_{S}$ by the principal eigenvector of the above matrix.
  • Refine this estimate using the remaining information $\boldsymbol{y}_{t,-}$.
We point out that Proposition 3.1 remains true in the regime $\sqrt{n}\ll k\ll n$, and hence we observe a gap between $t_{\mbox{\tiny\rm alg}}(n,k)$ and $t_{\mbox{\tiny\rm Bayes}}(n,k)$ in this regime as well. However, we are not aware of matching hardness results in this regime.
@@ 4 Reduction of estimation to diffusion-based sampling
In order to further support and clarify our main conclusions, we present two simple results showing that estimation in the Gaussian noise model is reducible to diffusion sampling. Informally, if diffusion sampling can be performed in polynomial time with sufficient accuracy, then we can sample in polynomial time from the posterior distribution of $\boldsymbol{x}\sim\mu$ given noisy information $\boldsymbol{y}=\boldsymbol{x}+\sigma\boldsymbol{g}$ (where $\boldsymbol{g}$ is standard Gaussian noise independent of $\boldsymbol{x}$).
Here we consider a general sampling problem for $\boldsymbol{x}\sim\mu$ a probability distribution in ${\mathbb{R}}^{d}$. For the sake of simplicity, we will assume $\mu$ to be supported on the unit sphere ${\mathbb{S}}^{d-1}=\{\boldsymbol{x}:\|\boldsymbol{x}\|=1\}$. Our assumptions will be about the accuracy of the denoising diffusion process. We denote by ${\rm P}_{\boldsymbol{y}}^{T}$ the law of $(\boldsymbol{y}_{t})_{0\leq t\leq T}$ where $\boldsymbol{y}_{t}$ given by Eq. (1) and by ${\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}$ the law of $(\hat{\boldsymbol{y}}_{t})_{0\leq t\leq T}$, which is the discretized diffusion trajectory defined in (4) (interpolated linearly outside ${\mathbb{N}}\cdot\Delta$).
@@
where $\lfloor t\rfloor_{\Delta}:=\max\{s\in{\mathbb{N}}\cdot\Delta:s\leq t\}$. It is easy to see that this definition interpolates $(\hat{\boldsymbol{y}}_{t}:t\in{\mathbb{N}}\cdot\Delta)$ defined in (4)
We state first an easier form of the reduction.
@@
where $\varepsilon_{0}(\theta):=\mathbb{E}\|{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}-{\mathsf{N}}(\boldsymbol{0},(\theta d)^{-1}\boldsymbol{I}_{d})*{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}\|_{\mbox{\tiny\rm TV}}$ is the expected TV distance between ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ and the convolution of ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ with a Gaussian with variance $1/(\theta d)$. As a consequence, there exists a randomized algorithm $\hat{\boldsymbol{m}}_{+}$ with complexity $(N\chi\cdot T/\Delta)$ that approximates the posterior expectation, namely
$\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{+}(\boldsymbol{y})-\mathbb{E}(\boldsymbol{x}|\boldsymbol{y})\|^{2}\big{\}}\leq 2\overline{\varepsilon}+\frac{2}{N}\,.$ (17)
Proof.
Finally, Eq. (17) follows by generating $N$ i.i.d. copies $\hat{\boldsymbol{x}}_{1},\dots,\hat{\boldsymbol{x}}_{N}$ using the above procedure, and letting $\hat{\boldsymbol{m}}(\boldsymbol{y})$ be their empirical average. ∎
The next statement makes a weaker assumption on the accuracy of the diffusion sampler (transportation instead of KL distance), but in exclnage assumes the approximate drift $\hat{\boldsymbol{m}}$ to be Lipschitz. We note that ${\rm Lip}(\boldsymbol{m}(\,\cdot\,,t))=\sup_{\boldsymbol{y}}\|{\rm Cov}(\boldsymbol{x}|\boldsymbol{y}_{t}=\boldsymbol{y})\|_{{\mbox{\rm\tiny op}}}$, and the latter is of $O(1/d)$ (for instance) if the coordinates of $\boldsymbol{x}$ are weakly dependent under the posterior.
@@
Assume that $\hat{\boldsymbol{m}}(\,\cdot\,,\,\cdot\,)$ has computational complexity $\chi$ and satisfies the following:
  1. $(a)$
    For every $t\geq 1/\sigma^{2}$, $\boldsymbol{y}\mapsto\hat{\boldsymbol{m}}(\boldsymbol{y},t)$ is $L/d$-Lipschitz.
  2. $(b)$
    There is a stepsize $\Delta$ such that $W_{1}({\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta},{\rm P}_{\boldsymbol{y}}^{T})\leq\varepsilon$ for any $T\leq\theta d$.

### v2

@@ 3 Main results
We will state two separate results for the probability distribution $\mu=\mu_{n,k}$ described in the introduction, depending on the scaling of $k$ with $n$: in Section 3.1 we assume $\sqrt{n}\ll k\ll n$; while in Section 3.3 we assume $k\ll\sqrt{n}$. In fact, the nature of the problem changes at the threshold $k\asymp\sqrt{n}$.
@@
It is expected that no polynomial-time algorithm can estimate $\boldsymbol{x}$ significantly better than the null estimator $\hat{\boldsymbol{m}}_{\mbox{\tiny\rm null}}=\boldsymbol{0}$ for $\underline{k}_{n}\leq k\ll n$, with $\underline{k}_{n}\ll n$ a suitable sequence (see Conjecture 3.2).
Since $\|\boldsymbol{x}\|_{F}=1$ for $\boldsymbol{x}\sim\mu$, the Bayes denoiser $\boldsymbol{m}(\boldsymbol{y},t)=\boldsymbol{m}(\boldsymbol{y})$ does not depend on $t$.
@@ 3.1 Moderately sparse regime: $\sqrt{n}\ll k\ll n$
Given an arbitrary polynomial time computable drift $\hat{\boldsymbol{m}}_{0}$, we construct a different poly-time drift $\hat{\boldsymbol{m}}$, with nearly equal score matching objective and yet incorrect sampling. In the next subsection we state a related result that establishes incorrect sampling from an explicit denoiser with near-optimal denoising properties.
@@
In order to connect this theorem with the discussion in the introduction, we recall two facts from the literature on submatrix estimation: $(i)$ The ideal posterior mean estimator $\boldsymbol{m}(\boldsymbol{y}_{t})$ achieves small mean square error in a large interval below $t_{\mbox{\tiny\rm alg}}$ (Proposition 3.1); $(ii)$ No polynomial-time estimator is expected to perform substantially better than the null estimator below $t_{\mbox{\tiny\rm alg}}$ (Conjecture 3.2).
Regarding $(i)$, we state a characterization of the Bayes optimal error. The proof is essentially the same (indeed simpler) as the main result in Butucea et al. (2015), which considers the case of asymmetric matrices. An alternative proof (for $k\leq n^{a}$, $a<5/6$) also follows from Barbier et al. (2020).
Proposition 3.1 (Modification of Butucea et al. (2015)).
Let $\boldsymbol{m}(\,\cdot\,)$ be the posterior mean estimator in Eq. (8). Assume $1\ll k\ll n$, and define $t_{\mbox{\tiny\rm Bayes}}(n,k):=2k\log(n/k)$. Then, for any $\delta>0$, we have
@@
Several groups have proved hardness results for estimating $\boldsymbol{u}$ in the model considered here, for $t$ sufficiently smaller than the spectral threshold $n/2$. In particular Brennan et al. (2018) establishes hardness conditional on hardness on the planted clique problem for $t\leq n^{1-\varepsilon}$, $\varepsilon>0$, Hopkins et al. (2017) proved failure of sum-of-squares relaxations also for $t\leq n^{1-\varepsilon}$, Schramm and Wein (2022) proved hardness for low-degree polynomials for the related hidden submatrix problem for $t\leq cn$, $c>0$ a sufficiently small constant. Finally, Kunisky et al. (2019) prove hardness for low-degree polynomials algorithms when $t\leq(1-\varepsilon)n/2$, albeit their results require $k$ to be of order $n$.
@@
It makes sense to assume that $\|\hat{\boldsymbol{m}}_{0}(\cdot,\cdot)\|_{F}\leq 1$. Since $\boldsymbol{x}\sim\mu_{n,k}$ and $\text{supp}(\mu_{n,k})\subset\{\boldsymbol{a}:\|\boldsymbol{a}\|_{F}\leq 1\}$ and the latter is a convex set, projecting any estimator $\hat{\boldsymbol{m}}_{0}$ onto this set yields a smaller MSE.
@@ 3.2 Moderately sparse regime: A concrete example
7:  Compute $\hat{S}$ by $\hat{S}:=\Big{\{}i\in[n]:\;|v_{t,i}|\geq\frac{\varepsilon}{\sqrt{k}}\Big{\}}$
@@
    $\phantom{A}\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C\,e^{-cn/k}\,.$
@@ 3.3 Very sparse regime: $k\ll\sqrt{n}$
3:  Let $\boldsymbol{g}_{t}\sim{\mathsf{N}}(0,t{\boldsymbol{I}}_{n\times n})$ and compute $\boldsymbol{y}_{t,+}:=\boldsymbol{y}_{t}+\sqrt{\varepsilon}\boldsymbol{g}_{t}$, $\boldsymbol{y}_{t,-}:=\boldsymbol{y}_{t}-\boldsymbol{g}_{t}/\sqrt{\varepsilon}$
@@
    $\phantom{A}\inf_{t\leq(1-\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}=1-o_{n}(1)\,,\;\;\;\sup_{t\geq(1+\delta)t_{\mbox{\tiny\rm alg}}}\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}(\boldsymbol{y}_{t},t)-\boldsymbol{x}\|^{2}\big{\}}\leq C(\delta,D)n^{-D}\,.$
@@
The pseudocode for the estimator $\hat{\boldsymbol{m}}(\,\cdot\,)$ that is constructed in the above is given as Algorithm 2. This is based on a standard approach in the literature Deshpande and Montanari (2016); Cai et al. (2017), with some modifications to allow for its analysis in the diffusion setting. The main steps are as follows: $(1)$ Perform Gaussian data splitting of $\boldsymbol{y}_{t}$ into $\boldsymbol{y}_{t,+}$, $\boldsymbol{y}_{t,-}$, see Line 3 of lgorithm 2, with most of the information preserved in $\boldsymbol{y}_{t,+}$. $(2)$ Use entrywise soft thresholding $\eta_{s}(x)=(|x|-s)_{+}\operatorname{sign}(x)$ to reduce the noise in the symmetrized version of $\boldsymbol{y}_{t,+}$. $(3)$ Compute a first estimate of the latent vector $\boldsymbol{1}_{S}$ by the principal eigenvector of the above matrix. $(4)$ Refine this estimate using the remaining information $\boldsymbol{y}_{t,-}$.
We point out that Proposition 3.1 remains true in the regime $\sqrt{n}\ll k\ll n$, and hence we observe a gap between $t_{\mbox{\tiny\rm alg}}(n,k)$ and $t_{\mbox{\tiny\rm Bayes}}(n,k)$ in this regime as well.
@@ 4 Reduction of estimation to diffusion-based sampling
In order to clarify our conclusions, we prove a general reduction: if diffusion sampling can be performed in polynomial time with sufficient accuracy, then we can sample in polynomial time from the posterior distribution of $\boldsymbol{x}\sim\mu$ given noisy information $\boldsymbol{y}=\boldsymbol{x}+\sigma\boldsymbol{g}$ (where $\boldsymbol{g}$ is standard Gaussian noise independent of $\boldsymbol{x}$). By repeating this sampling, we can also perform estimation.
Here we consider a general sampling problem for $\boldsymbol{x}\sim\mu$ a probability distribution in ${\mathbb{R}}^{d}$. For the sake of simplicity, we will assume $\mu$ to be supported on the unit sphere ${\mathbb{S}}^{d-1}=\{\boldsymbol{x}:\|\boldsymbol{x}\|=1\}$. We denote by ${\rm P}_{\boldsymbol{y}}^{T}$ the law of $(\boldsymbol{y}_{t})_{0\leq t\leq T}$ where $\boldsymbol{y}_{t}$ given by Eq. (1) and by ${\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta}$ the law of $(\hat{\boldsymbol{y}}_{t})_{0\leq t\leq T}$, which is the discretized diffusion trajectory defined in (4) (interpolated linearly outside ${\mathbb{N}}\cdot\Delta$).
@@
where $\lfloor t\rfloor_{\Delta}:=\max\{s\in{\mathbb{N}}\cdot\Delta:s\leq t\}$. This definition interpolates $(\hat{\boldsymbol{y}}_{t}:t\in{\mathbb{N}}\cdot\Delta)$ defined in (4).
@@
where $\varepsilon_{0}(\theta):=\mathbb{E}\|{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}-{\mathsf{N}}(\boldsymbol{0},(\theta d)^{-1}\boldsymbol{I}_{d})*{\rm P}_{\boldsymbol{x}|\boldsymbol{y}}\|_{\mbox{\tiny\rm TV}}$ is the expected TV distance between ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ and the convolution of ${\rm P}_{\boldsymbol{x}|\boldsymbol{y}}$ with a Gaussian with variance $1/(\theta d)$. As a consequence, there exists a randomized algorithm $\hat{\boldsymbol{m}}_{+}$ with complexity $(N\chi\cdot T/\Delta)$ that approximates the posterior expectation:
$\mathbb{E}\big{\{}\|\hat{\boldsymbol{m}}_{+}(\boldsymbol{y})-\mathbb{E}(\boldsymbol{x}|\boldsymbol{y})\|^{2}\big{\}}\leq 2\overline{\varepsilon}+2N^{-1}\,.$ (14)
The next statement makes a weaker assumption on the accuracy of the diffusion sampler (transportation instead of KL distance), but in exchnage assumes the approximate drift $\hat{\boldsymbol{m}}$ to be Lipschitz. We note that ${\rm Lip}(\boldsymbol{m}(\,\cdot\,,t))=\sup_{\boldsymbol{y}}\|{\rm Cov}(\boldsymbol{x}|\boldsymbol{y}_{t}=\boldsymbol{y})\|_{{\mbox{\rm\tiny op}}}$, and the latter is of $O(1/d)$ (for instance) if the coordinates of $\boldsymbol{x}$ are weakly dependent under the posterior.
@@
Assume that $\hat{\boldsymbol{m}}(\,\cdot\,,\,\cdot\,)$ has computational complexity $\chi$ and satisfies the following: $(a)$ For every $t\geq 1/\sigma^{2}$, $\boldsymbol{y}\mapsto\hat{\boldsymbol{m}}(\boldsymbol{y},t)$ is $L/d$-Lipschitz. $(b)$ There is a stepsize $\Delta$ such that $W_{1}({\rm P}_{\hat{\boldsymbol{y}}}^{T,\Delta},{\rm P}_{\boldsymbol{y}}^{T})\leq\varepsilon$ for any $T\leq\theta d$.